Grade 12th passMechanics

Required energy to move an object of mass m from an orbit of radius R to radius 2R ( M is mass of planet)
  1. 2GMm/R
  2. GMm/2R
  3. GMm/4R
  4. GMm/8R

Aman
8 Years agoGrade 12th pass

1 Answer

Arun
8 Years ago
 
Energy required = (Potential energy of the Earth – mass system when mass is at distance 3R) – (Potential energy of the Earth-mass system when mass is at distance 2R)
= - GM/3R – (-GM/2R) = - GM/3R + GM/2R
= -2GMm+3GMm / 6R = GMm/6R 
Energy required = (Potential energy of the Earth – mass system when mass is at distance 2R) – (Potential energy of the Earth-mass system when mass is at distance R)
= - GM/2R – (-GM/R) = - GM/2R + GM/R
= -GMm+2GMm / 2R = GMm/2R
 
hence option B is correct